Fermi National Laboratory


Accelerator Update for April 18, 2005 - April 20, 2005

Monday April 18
The day shift began with Operations monitoring store 4100 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with the MiniBooNE and the Meson MCenter experiments taking beam.

At 9:08 AM, a Booster power supply (EAPS) tripped off and would not reset. Operators investigated and turned the supply on locally.

At 9:52 AM, Operations reported an emittance blow up in Pbar.

At 10:09 AM, a Main Injector (MI) expert asked Operations to investigate a RF cavity phase error. Operators verified the problem and then contacted RF experts.

Operations terminated store 4100 at 11:04 AM.

Operators prepared the Collider Detector at Fermilab (CDF) collision hall for access at 11:21 AM.

Cryo system technicians began maintenance on wet engines at Tevatron (TeV) sectors F1 and A1 at 12:10 PM. They completed their work an hour later.

CDF came out of access at 12:46 PM.

Operators dumped the Pbar stack at 12:58 PM so techs could work on an Accumulator quadrupole power supply (A:LQ). Experts will also access the pre-vault tunnel to check on a toroid (TOR105).

Operations put the TeV through a dry squeeze at 1:10 PM.

Operations put the TeV into shot setup at 4:21 PM.

At 5:26 PM, Operations reported problems with TLCK. An operator called in Controls experts. The experts eventually replaced the CAMAC link repeaters at E3 and E4.

Operations resumed shot setup at 7:55 PM.

A Pbar power supply expert began work on A:LQ at 9:55 PM.

At 9:56 PM, a problem with the Booster beam required an expert to be called in.

Operators prepared the Pbar pre-vault tunnel for access at 10:11 PM.

Operations established combination store 4104 at 10:17 PM with an initial luminosity of 91.7E30.

Operations established beam to MCenter at 10:24 PM.

Operations established beam to MiniBooNE at 10:45 PM.

At 11:08 PM, a Booster RF station (BRF16) tripped off. It reset okay.

Operations reported at 11:10 PM that work on A:LQ was finished.

At 11:47 PM, Operations tried to resume stacking, but another problem with A:LQ appeared.

Tuesday April 19
The midnight shift began with Operations monitoring store 4104, with operators talking with an expert about A:LQ problems, and with no beam to MiniBooNE or MCenter due to a Booster notcher problem.

Operations reported at 12:09 AM that they swapped a number of components for A:LQ, per expert instructions, with another power supply. They had it working by 1 AM.

Pbar resumed stacking 1:11 AM.

At 8:02 AM, Operations reported that a Booster vacuum burst had tripped off several vacuum pumps. They all reset okay.

Operations began stashing antiprotons into the Recycler at 8:20 AM. The stash was completed by 9:16 AM.

Pbar resumed stacking at 9:22 PM.

MCenter resumed taking beam at 9:49 AM.

At 11:27 AM, the Booster notcher tripped off again. Experts investigated and had it back up in ten minutes.

Operations began stashing antiprotons into the Recycler at 1:31 PM. This transfer ended an hour later and MCenter resumed taking beam.

Pbar resumed stacking at 3:27 PM.

At 6:54 PM, Operations reported that an Accumulator Septa (A:ISEP2V) tripped off and wouldn't reset. Operations called in an expert.

The Pbar power supply expert had A:ISEP2V back on line and Pbar stacking by 7:55 PM.

At 8:17 PM, Operations reported that stacking had stopped. Operators investigated and found a problem with A:ISEP2V. They called in an expert.

The Pbar expert reported at 11:32 PM that he had replaced some bad SCRs, but there was still a problem with A:ISEP2V.

Wednesday April 20
The midnight shift began with Operations monitoring store 4104 and the stash, with stacking off due to A:ISEP2V repairs, and with MiniBooNE and MCenter taking beam.

The Pbar power supply expert reported 2:51 AM that he replaced all the SCRs for A:ISEP2V and that it was firing normally. Pbar resumed stacking.

At 7:40 AM, BRF16 tripped off. It reset okay.

The Plan for Wednesday
The plan for today is to terminate the store around 8 AM and then put the TeV back into shot setup. Also, at the end of shift A:ISEP2V started misfiring again, which halted stacking; experts investigated.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/20/2005   email Fermilab